# An Analysis of Action-Value Temporal-Difference Methods That Learn State Values

### v1

@@ 2 Background
Most of the algorithms considered in this paper are off-policy methods. They learn an action-value function $Q(s,a)$ that estimates either $q_{\pi}(s,a)$ or $q_{*}(s,a)$ using samples obtained from executing policy $b$ in the environment. For example, Expected Sarsa (John, 1994) is an off-policy TD method that uses explicit knowledge of the target policy $\pi$ to approximate the solution to Eq. 1 from sampled data:
@@
where $\alpha\in(0,1]$ is the step size of the update. We revisit Expected Sarsa as an important baseline several times in this paper because it generalizes a number of fundamental bootstrapping algorithms (van Hasselt, 2011). For instance, when the target policy is greedy with respect to $Q$, the expectation in Eq. 2 becomes equivalent to $\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime})$, yielding the Q-learning algorithm:
@@
which will be useful for the control case in Section 3.2.
@@ 2.1 QV-learning
The state-value function is related to the action-value function by $v_{\pi}(s)=\sum_{a\in\mathcal{A}}\pi(a|s)q_{\pi}(s,a)$. This implies that an approximation of $v_{\pi}(s^{\prime})$ can be substituted in Eq. 1 to avoid the summation over actions. QV-learning (Wiering, 2005) is an on-policy TD algorithm based on this concept. The idea is to use TD(0) to independently learn $V(s)\approx v_{b}(s)$, while concurrently bootstrapping from these estimated state values to learn $Q(s,a)\approx q_{b}(s,a)$. The specific value-function updates for QV-learning become
@@
Wiering (2005, Sec. 2) is explicit that $Q$ is always updated before $V$ in each iteration. Additionally, it is commonly assumed that both updates share the same step size, $\alpha$, as is shown above. We analyze this algorithm in Section 3.1.
The main variant of QV-learning for off-policy control is QVMAX (Wiering & van Hasselt, 2009), which essentially substitutes the max operator from Q-learning into Eq. 6 in an attempt to induce convergence to $q_{*}$ instead of $q_{b}$:
@@
The update to $Q(S_{t},A_{t})$ remains the same as Eq. 5 and is still conducted prior to Eq. 7 at each time step. We analyze this algorithm in Section 3.2.
@@ 2.2 Dueling Q-learning
where the last term is subtracted to make the solution unique (we further discuss this point in Section 4.2). The updates to $V(s)$ and $\operatorname{Adv}(s,a)$ are then derived implicitly from Q-learning by reframing its update as stochastic semi-gradient descent. Let ${\bm{\theta}}$ be the vector of all learnable value-function parameters involved in Eq. 8. We therefore write $Q(s,a;{\bm{\theta}})$ for the action-value estimate of state-action pair $(s,a)$. Additionally, let ${\bm{\theta}}^{-}$ be the target parameters for bootstrapping (Mnih et al., 2015). The Q-learning update in Eq. 3 is rewritten as the following squared-error loss minimization:
@@
Applying the chain rule to $\mathop{\nabla_{\!{\bm{\theta}}}}Q(S_{t},A_{t};{\bm{\theta}})$ in Eq. 9 according to the value decomposition in Eq. 8 yields the final dueling update. In practice, the decomposition in Eq. 8 is represented by a branching and merging neural network, and the chain rule is handled by automatic differentiation. In Section 4.1, we instead apply the chain rule to Eq. 8 manually, assuming tabular value functions, to derive a Dueling Q-learning update without function approximation. This allows us to gain further insight into this fundamental algorithm beyond the deep RL setting.
@@ 3.1 On-Policy Prediction
To test this, we design a parametric MDP experiment where the environment has $4$ states and $\left\lvert\mathcal{A}\right\rvert$ actions. Taking any action in any state triggers a transition to a random state (possibly the same one) with equal probability. The agents’ behavior policy is uniform random in every state. Agents receive a reward of $+1$ whenever the first action, $a_{0}$, is taken, and a reward of 00 otherwise.
@@
Because the MDP dynamics are state-invariant, $v_{b}(s)$ is constant for all $s\in\mathcal{S}$. Solving the state-value Bellman equation, Eq. 4, yields $v_{b}(s)=1\mathbin{/}\big{(}(1-\gamma)\,\left\lvert\mathcal{A}\right\rvert\big{)}$. It follows that $q_{b}(s,a_{i})={\bm{1}}_{\{i=0\}}+\gamma\mathbin{/}\big{(}(1-\gamma)\,\left\lvert\mathcal{A}\right\rvert\big{)}$ for all $a_{i}\in\mathcal{A}$, where $\mathbf{1}$ is the indicator function. We set $\gamma=0.99$.
@@
We evaluate the agents’ performance by measuring the root mean square (RMS) prediction error, $\|q_{b}-Q\|_{2}$, as a function of the number of environment interactions. We normalize the errors by expressing them as a percentage of the initial RMS error, $\|q_{b}-Q_{0}\|_{2}$, where $Q_{0}$ is initialized with zeros. We compare Expected Sarsa and QV-learning, training both agents for 20,000 time steps. In Figure 1 (left), we plot the learning curves for the case where $\left\lvert\mathcal{A}\right\rvert=18$, the largest instantiation of our action set that we consider here. We then average the results over 100 independent trials, with shading indicating 95% confidence intervals. The agents’ step sizes are chosen from a natural-logarithmic grid search of 61 values over the interval $(0,1]$, similarly to the random walk experiment by Sutton & Barto (2018, Sec. 12.1). The selection criterion is to minimize the area under the curve (AUC), which we normalized as a percentage. Specifically, an AUC of 100% indicates that the initial error did not change at all, whereas a lower percentage indicates faster average progress towards the fixed point. In Figure 1 (center), we plot the corresponding AUCs for each step size tested by the sweep. The dashed horizontal line corresponds to the smallest AUC achieved by each agent.
@@
Figure 1: On-policy prediction performance of QV-learning compared to Expected Sarsa in the 4-state MDP. (Left) Root mean square (RMS) prediction error versus the number of learning steps, after optimizing the step size, $\alpha\in(0,1]$. (Center) Area under the learning curve (AUC), normalized as a percentage, for each step size. The dashed lines correspond to the AUCs for the learning curves in the left subplot. (Right) The smallest AUC obtained across all step sizes, as the number of MDP actions is increased. Results are averaged over 100 independent trials. Shading represents 95% confidence intervals.
@@
To investigate the scalability of these methods, we repeat this experiment setup for ${\left\lvert\mathcal{A}\right\rvert\in\{2,6,10,14,18\}}$. We then plot the best AUC achieved in each problem instance in Figure 1 (right). The slope of the resulting lines roughly correspond to the scalability of the algorithms. We note that QV-learning’s line is more horizontal, indicating better scaling to large action-space cardinalities, as we hypothesized earlier. However, the trade-off for this improved sample efficiency is a much noisier update, which can be seen in Figure 1 (left); if we were to train for longer, Expected Sarsa would eventually obtain a more accurate solution. This indicates that QV-learning is perhaps best suited for cases where learning a satisfactory—but possibly imperfect—value function is desired within a small number of samples.
@@
These results do not, however, suggest that QV-learning fails to converge to the correct pair of on-policy value functions, $q_{b}$ and $v_{b}$. The noisy behavior observed in Figure 1 (left) is primarily due to the fact that the step size is not annealed. The following theorem states that, in expectation, the joint update of QV-learning is a contraction mapping when we represent the two value functions as a concatenated vector of size $\left\lvert\mathcal{S}\right\rvert+\left\lvert\mathcal{S}\times\mathcal{A}\right\rvert$.
@@
See Section A.1. ∎
The significance of this result is that it shows that both of QV-learning’s value functions make progress (on average) towards their respective fixed points, without requiring $V$ to converge first. This is in contrast to less formal convergence arguments for QV-learning, which may rely on the assumption that TD(0) would first converge to $v_{b}$ based on well-established convergence results, and then the update rule in Eq. 5 would be able to bootstrap from it to extract $q_{b}$. The issue with this two-timescale approach is that $V$ may never exactly equal $v_{b}$ after any finite amount of time, meaning that $Q$ would still incur some bootstrapping bias—a caveat that is directly addressed by considering the joint operator space, as we have done here.
@@ 3.2 Off-Policy Control
One consequence of this change is that now there is reciprocal bootstrapping between $Q$ and $V$. Previously, with QV-learning, we had only a unidirectional information flow: $V$ bootstrapped from itself, and $Q$ bootstrapped from $V$. Hence, $Q$ could not corrupt the state values in any way. We illustrate this distinction in Figure 2. Although reciprocal bootstrapping is not necessarily an undesirable property, it is worth noting that a fundamental characteristic of the algorithm has changed.
Another, more serious consequence of this change is that the update now suffers from off-policy bias. Wiering & van Hasselt (2009) hypothesized that the use of the max operator in Eq. 7 would induce convergence to $q_{*}$, just as it does in Q-learning. Unfortunately, this is not true, as we show in the next proposition.
@@
$\begin{bmatrix}{\bm{q}}_{*}\\ {\bm{v}}_{*}\end{bmatrix}$ is generally not a fixed point of QVMAX.
@@
See Section A.2. ∎
The update fails to correct the distribution of the observed reward, $R_{t+1}$, which is collected by the behavior policy. This is because Eq. 7 is not conditioned on the state-action pair, $(S_{t},A_{t})$, as is normally the case for Q-learning, but is instead conditioned only on the state, $S_{t}$. The freedom in the choice of $A_{t}$ means the expected reward observed from state $S_{t}$ now depends on behavior policy $b$.
To eliminate the bias, the update cannot depend explicitly on the sampled reward, $R_{t+1}$. We therefore propose to modify Eq. 7 in the QVMAX updates such that we have
@@
See Section A.3. ∎
A peculiarity of the updates in Eqs. 5 and 10 is that they closely resemble Q-learning in Eq. 3. In fact, if we set $\alpha=1$, then Eq. 10 simply becomes a table overwrite: $V(S_{t})\leftarrow\max_{a\in\mathcal{A}}Q(S_{t},a)$. This would make the algorithm almost the same as Q-learning, but with a delay introduced before bootstrapping. When $\alpha<1$, then Eq. 10 has an extra smoothing effect, mixing together stale estimates of the maximum action value in each state and exacerbating this delay.
We slightly modify the prediction experiment from Section 3.1 to demonstrate the bias of QVMAX. This time, we target a greedy policy with respect to the current action-value function, making Expected Sarsa equivalent to Q-learning. We therefore compare Q-learning, QVMAX, and BC-QVMAX. The optimal policy is to select $a_{0}$ unconditionally, making $q_{*}(s,a_{0})=1\mathbin{/}(1-\gamma)$. It follows that $q_{*}(s,a_{i})=\gamma\mathbin{/}(1-\gamma)$ for $i\neq 0$.
All experiment and plotting procedures remain the same as before, except that now the RMS error is defined in terms of $q_{*}$: i.e., $\|q_{*}-Q\|_{2}$. We plot the results in Figure 3. As can be seen in Figure 3 (left), and as our theory predicted, QVMAX asymptotes substantially above the zero-error mark due to its biased update. In contrast, BC-QVMAX achieves a much lower error. However, as we also discussed above, the smoothing effect of BC-QVMAX makes it similar to, but rather slower than, Q-learning. The results in Figure 3 (left) are identical only because $\alpha=1$ happens to be the best step size for both methods, which makes them nearly equivalent (except for the delay mentioned previously). However, for all $\alpha<1$, we can see from Figure 3 (center) that BC-QVMAX slightly lags behind Q-learning.
@@
Figure 3: Off-policy control performance of BC-QVMAX compared to QVMAX and Q-learning in the 4-state MDP. Experiment setup is the same as that of Figure 1. (Note that Q-learning is sometimes eclipsed by BC-QVMAX.)
@@ 4.1 Dueling Q-learning
To obtain a tabular Dueling Q-learning update, we evaluate the chain rule in Eq. 9, as discussed earlier in Section 2.2. Q-learning’s TD error is $\delta_{t}=R_{t+1}+\gamma\max_{a^{\prime}\in\mathcal{A}}Q(S_{t+1},a^{\prime})-Q(S_{t},A_{t})$. Because $Q(s,a)$ is defined according to Eq. 8, the chain rule preserves the error, $\delta_{t}$, inside the quadratic term and then scales it in the following manner:

### v2

@@ 2 Background
Most of the algorithms considered in this paper are off-policy methods. They learn an action-value function $Q(s,a)$ that estimates either $q_{\pi}(s,a)$ or $q_{*}(s,a)$ using samples obtained from executing policy $b$ in the environment. For example, Expected Sarsa (John, 1994) is an off-policy TD method that uses explicit knowledge of the target policy $\pi$ to approximate the solution to Eq.˜1 from sampled data:
@@
where $\alpha\in(0,1]$ is the step size of the update. We revisit Expected Sarsa as an important baseline several times in this paper because it generalizes a number of fundamental bootstrapping algorithms (van Hasselt, 2011). For instance, when the target policy is greedy with respect to $Q$, the expectation in Eq.˜2 becomes equivalent to $\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime})$, yielding the Q-learning algorithm:
@@
which will be useful for the control case in Section˜3.2.
@@ 2.1 QV-learning
The state-value function is related to the action-value function by $v_{\pi}(s)=\sum_{a\in\mathcal{A}}\pi(a|s)q_{\pi}(s,a)$. This implies that an approximation of $v_{\pi}(s^{\prime})$ can be substituted in Eq.˜1 to avoid the summation over actions. QV-learning (Wiering, 2005) is an on-policy TD algorithm based on this concept. The idea is to use TD(0) to independently learn $V(s)\approx v_{b}(s)$, while concurrently bootstrapping from these estimated state values to learn $Q(s,a)\approx q_{b}(s,a)$. The specific value-function updates for QV-learning become
@@
Wiering (2005, Sec. 2) is explicit that $Q$ is always updated before $V$ in each iteration. Additionally, it is commonly assumed that both updates share the same step size, $\alpha$, as is shown above. We analyze this algorithm in Section˜3.1.
The main variant of QV-learning for off-policy control is QVMAX (Wiering & van Hasselt, 2009), which essentially substitutes the max operator from Q-learning into Eq.˜6 in an attempt to induce convergence to $q_{*}$ instead of $q_{b}$:
@@
The update to $Q(S_{t},A_{t})$ remains the same as Eq.˜5 and is still conducted prior to Eq.˜7 at each time step. We analyze this algorithm in Section˜3.2.
@@ 2.2 Dueling Q-learning
where the last term is subtracted to make the solution unique (we further discuss this point in Section˜4.2). The updates to $V(s)$ and $\operatorname{Adv}(s,a)$ are then derived implicitly from Q-learning by reframing its update as stochastic semi-gradient descent. Let $\bm{\theta}$} be the vector of all learnable value-function parameters involved in Eq.˜8. We therefore write $Q(s,a;{\bm{\theta}})$ for the action-value estimate of state-action pair $(s,a)$. Additionally, let $\bm{\theta}$$^{-}$ be the target parameters for bootstrapping (Mnih et al., 2015). The Q-learning update in Eq.˜3 is rewritten as the following squared-error loss minimization:
@@
Applying the chain rule to $\mathop{\nabla_{\!{\bm{\theta}}}}Q(S_{t},A_{t};{\bm{\theta}})$ in Eq.˜9 according to the value decomposition in Eq.˜8 yields the final dueling update. In practice, the decomposition in Eq.˜8 is represented by a branching and merging neural network, and the chain rule is handled by automatic differentiation. In Section˜4.1, we instead apply the chain rule to Eq.˜8 manually, assuming tabular value functions, to derive a Dueling Q-learning update without function approximation. This allows us to gain further insight into this fundamental algorithm beyond the deep RL setting.
@@ 3.1 On-Policy Prediction
To test this, we design a parametric MDP experiment where the environment has $4$ states and $\left\lvert\mathcal{A}\right\rvert$ actions. Taking any action in any state triggers a transition to a random state (possibly the same one) with equal probability. The agents’ behavior policy is uniform random in every state. Agents receive a reward of +1+1 whenever the first action, $a_{0}$, is taken, and a reward of 0 otherwise.
@@
Because the MDP dynamics are state-invariant, $v_{b}(s)$ is constant for all $s\in\mathcal{S}$. Solving the state-value Bellman equation, Eq.˜4, yields $v_{b}(s)=1\mathbin{/}\big{(}(1-\gamma)\,\left\lvert\mathcal{A}\right\rvert\big{)}$. It follows that $q_{b}(s,a_{i})={\bm{1}}_{\{i=0\}}+\gamma\mathbin{/}\big{(}(1-\gamma)\,\left\lvert\mathcal{A}\right\rvert\big{)}$ for all $a_{i}\in\mathcal{A}$, where $\mathbf{1}$ is the indicator function. We set $\gamma=0.99$.
@@
We evaluate the agents’ performance by measuring the root mean square (RMS) prediction error, $\|q_{b}-Q\|_{2}$, as a function of the number of environment interactions. We normalize the errors by expressing them as a percentage of the initial RMS error, $\|q_{b}-Q_{0}\|_{2}$, where $Q_{0}$ is initialized with zeros. We compare Expected Sarsa and QV-learning, training both agents for 20,000 time steps. In Figure˜1 (left), we plot the learning curves for the case where $\left\lvert\mathcal{A}\right\rvert=18$, the largest instantiation of our action set that we consider here. We then average the results over 100 independent trials, with shading indicating 95% confidence intervals. The agents’ step sizes are chosen from a natural-logarithmic grid search of 61 values over the interval $(0,1]$, similarly to the random walk experiment by Sutton & Barto (2018, Sec. 12.1). The selection criterion is to minimize the area under the curve (AUC), which we normalized as a percentage. Specifically, an AUC of 100% indicates that the initial error did not change at all, whereas a lower percentage indicates faster average progress towards the fixed point. In Figure˜1 (center), we plot the corresponding AUCs for each step size tested by the sweep. The dashed horizontal line corresponds to the smallest AUC achieved by each agent.
@@
Figure 1: On-policy prediction performance of QV-learning compared to Expected Sarsa in the 4-state MDP. (Left) Root mean square (RMS) prediction error versus the number of learning steps, after optimizing the step size, $\alpha\in(0,1]$. (Center) Area under the learning curve (AUC), normalized as a percentage, for each step size. The dashed lines correspond to the AUCs for the learning curves in the left subplot. (Right) The smallest AUC obtained across all step sizes, as the number of MDP actions is increased. Averaged across 100 trials; shading represents 95% confidence intervals.
@@
To investigate the scalability of these methods, we repeat this experiment setup for $\left\lvert\mathcal{A}\right\rvert\in\{2,6,10,14,18\}$}. We then plot the best AUC achieved in each problem instance in Figure˜1 (right). The slope of the resulting lines roughly correspond to the scalability of the algorithms. We note that QV-learning’s line is more horizontal, indicating better scaling to large action-space cardinalities, as we hypothesized earlier. However, the trade-off for this improved sample efficiency is a much noisier update, which can be seen in Figure˜1 (left); if we were to train for longer, Expected Sarsa would eventually obtain a more accurate solution. This indicates that QV-learning is perhaps best suited for cases where learning a satisfactory—but possibly imperfect—value function is desired within a small number of samples.
@@
These results do not, however, suggest that QV-learning fails to converge to the correct pair of on-policy value functions, $q_{b}$ and $v_{b}$. The noisy behavior observed in Figure˜1 (left) is primarily due to the fact that the step size is not annealed. The following theorem states that, in expectation, the joint update of QV-learning is a contraction mapping when we represent the two value functions as a concatenated vector of size $\left\lvert\mathcal{S}\right\rvert+\left\lvert\mathcal{S}\times\mathcal{A}\right\rvert$.
@@
See Section˜A.1. ∎
The significance of this result is that it shows that both of QV-learning’s value functions make progress (on average) towards their respective fixed points, without requiring $V$ to converge first. This is in contrast to less formal convergence arguments for QV-learning, which may rely on the assumption that TD(0) would first converge to $v_{b}$ based on well-established convergence results, and then the update rule in Eq.˜5 would be able to bootstrap from it to extract $q_{b}$. The issue with this two-timescale approach is that $V$ may never exactly equal $v_{b}$ after any finite amount of time, meaning that $Q$ would still incur some bootstrapping bias—a caveat that is directly addressed by considering the joint operator space, as we have done here.
@@ 3.2 Off-Policy Control
One consequence of this change is that now there is reciprocal bootstrapping between $Q$ and $V$. Previously, with QV-learning, we had only a unidirectional information flow: $V$ bootstrapped from itself, and $Q$ bootstrapped from $V$. Hence, $Q$ could not corrupt the state values in any way. We illustrate this distinction in Figure˜2. Although reciprocal bootstrapping is not necessarily an undesirable property, it is worth noting that a fundamental characteristic of the algorithm has changed.
A more serious consequence of this change is that the update now suffers from off-policy bias. Wiering & van Hasselt (2009) hypothesized that the max operator in Eq.˜7 would induce convergence to $q_{*}$, as it does in Q-learning. Unfortunately, this is not true, as we show in the next proposition.
@@
Unless the behavior policy, $b$, satisfies ${\mathbb{E}_{b}[R_{t+1}+\gamma\mathop{v_{*}(S_{t+1})}\mid S_{t}=s]}={\mathbb{E}_{\pi_{*}}[R_{t+1}+\gamma\mathop{v_{*}(S_{t+1})}\mid S_{t}=s]}$ for all $s\in\mathcal{S}$, then $\begin{bmatrix}{\bm{q}}_{*}\\ {\bm{v}}_{*}\end{bmatrix}$ is not a fixed point of QVMAX.
@@
See Section˜A.2. ∎
The update fails to correct the distribution of the observed reward, $R_{t+1}$, which is collected by the behavior policy. This is because Eq.˜7 is not conditioned on the state-action pair, $(S_{t},A_{t})$, as is normally the case for Q-learning, but is instead conditioned only on the state, $S_{t}$. The freedom in the choice of $A_{t}$ means the expected reward observed from state $S_{t}$ now depends on behavior policy $b$.
To eliminate the bias, the update cannot depend explicitly on the sampled reward, $R_{t+1}$. We therefore propose to modify Eq.˜7 in the QVMAX updates such that we have
@@
See Section˜A.3. ∎
A peculiarity of the updates in Eqs.˜5 and 10 is that they closely resemble Q-learning in Eq.˜3. In fact, if we set $\alpha=1$, then Eq.˜10 simply becomes a table overwrite: $V(S_{t})\leftarrow\max_{a\in\mathcal{A}}Q(S_{t},a)$. This would make the algorithm almost the same as Q-learning, but with a delay introduced before bootstrapping. When $\alpha<1$, then Eq.˜10 has an extra smoothing effect, mixing together stale estimates of the maximum action value in each state and exacerbating this delay.
We slightly modify the prediction experiment from Section˜3.1 to demonstrate the bias of QVMAX. This time, we target a greedy policy with respect to the current action-value function, making Expected Sarsa equivalent to Q-learning. We therefore compare Q-learning, QVMAX, and BC-QVMAX. The optimal policy is to select $a_{0}$ unconditionally, making $q_{*}(s,a_{0})=1\mathbin{/}(1-\gamma)$. It follows that $q_{*}(s,a_{i})=\gamma\mathbin{/}(1-\gamma)$ for $i\neq 0$.
All experiment and plotting procedures remain the same as before, except that now the RMS error is defined in terms of $q_{*}$: i.e., $\|q_{*}-Q\|_{2}$. We plot the results in Figure˜3. As can be seen in Figure˜3 (left), and as our theory predicted, QVMAX asymptotes substantially above the zero-error mark due to its biased update. In contrast, BC-QVMAX achieves a much lower error. However, as we also discussed above, the smoothing effect of BC-QVMAX makes it similar to, but rather slower than, Q-learning. The results in Figure˜3 (left) are identical only because $\alpha=1$ happens to be the best step size for both methods, which makes them nearly equivalent (except for the delay mentioned previously). However, for all $\alpha<1$, we can see from Figure˜3 (center) that BC-QVMAX slightly lags behind Q-learning.
@@
Figure 3: Off-policy control performance of BC-QVMAX compared to QVMAX and Q-learning in the 4-state MDP. Experiment setup is the same as that of Figure˜1. (Note that Q-learning is sometimes eclipsed by BC-QVMAX.)
@@ 4.1 Dueling Q-learning
To obtain a tabular Dueling Q-learning update, we evaluate the chain rule in Eq.˜9, as discussed earlier in Section˜2.2. Q-learning’s TD error is $\delta_{t}=R_{t+1}+\gamma\max_{a^{\prime}\in\mathcal{A}}Q(S_{t+1},a^{\prime})-Q(S_{t},A_{t})$. Because $Q(s,a)$ is defined according to Eq.˜8, the chain rule preserves the error, $\delta_{t}$, inside the quadratic term and then scales it in the following manner: